TEX Users Group

P.O. Box 869 Santa Barbara, California 93102 USA

Founded:  In 1980, in Providence, Rhode Island (USA).
Mission Statement: The TEX Users Group (TUG) provides leadership:

▷ to encourage and expand the use of TEX, , and related

systems ▷ to ensure the integrity and portability of TEX, , and related

systems ▷ to foster innovation in high-quality electronic document preparation

Officers:

President  ▷ Christina Thiele (cthiele@ccs.carleton.ca)
Vice-President  ▷ Michel Goossens (m.goossens@cern.ch)
Secretary  ▷ Peter Flynn (pflynn@curia.ucc.ie)
Treasurer  ▷ George Greenwade (bed gdg@shsu.edu)

Executive

Director:  ▷ Patricia Monohon
Membership:  ▷ ca. 2,300

Regular

Publications:  ▷ TUGboat (quarterly journal)

▷ TEX and TUG NEWS (quarterly newsletter)

▷ annual membership directory

Annual Meetings:  ▷ 1994: Santa Barbara, California (July 31–August 4, 1994)

▷ 1995: St. Petersburg Beach, Florida (July 24–28, 1995)

▷ 1996: Europe (site not yet selected)

Other Publications:

▷ TEXniques — manuals and guides to TEX and LaTEX, as well as TEX-related software such as PICTEX (graphics package) and edmac.sty (for critical editions; forthcoming).

“One Stop Shop”:  ▷ all TEX publications; software on disc and CD

▷ 20% discount for members

Courses:  ▷ from beginner to advanced levels, on TEX and related topics
Other Information:^{1} ▷ subdirectory on CTAN: /tex-archive/usergrps/tug

▷ directory of TEX consultants

▷ introduction to CTAN, the Comprehensive TEX Archive Network ▷ lists of TEX implementations for various platforms, other user groups

around the world, publishers accepting TEX files …

1 Some of these are still being developed.

Tel.: +1 (805) 963-1338  tug@tug.org  Fax: +1 (805) 963-8358

Source: /tex-archive/usergrps/tug/tug-info.tex on CTAN. Rev. 25 Nov. 94